A diffraction grating, ruled with 300 lines per mm, is illuminated with a white light source at normal incidence. (i) What is the angular separation, in the third-order spectrum, between the 400 nm and 600 nm lines? [5] (ii) Water (of refractive index 1.33) now fills the whole space between the grating and the screen? What is the angular separation, in the first-order spectrum, between the 400 nm and 600 nm lines? [5]
1
Expert's answer
2021-10-24T18:27:30-0400

(i)

"d= \frac{1}{N_p} = \frac{10^{-3}}{300} \\

N_p = 300 \times 10^3 \\

λ_1 = 400 \times 10^{-9} \;m \\

λ_2 = 600 \times 10^{-9} \;m \\

dsin θ_1 = nλ_1 \\

\frac{10^{-3}}{300} \times sin θ_1 = 3 \times 400 \times 10^{-9} \\

sin θ_1 = 12 \times 3 \times 10^{-2} \\

θ_1 = sin^{-1} (0.36) \\

θ_1=21.1 \;° \\

\frac{10^{-3}}{300} sin θ_2 = 3 \times λ_2 \\

\frac{10^{-3}}{300} sin θ_2 = 3 \times 600 \times 10^{-9} \\

θ_2 = sin^{-1} (0.54) \\

θ_2 = 32.68 \;° \\

Δ θ = θ_2- θ_1 \\

Δ θ = 32.68-21.1 = 11.58 \;°"

(ii)

"dsin θ^*_1 = \frac{mλ_1}{n} \\

n = 1.33 \\

m=3 \\

θ^*_1 = sin^{-1} (\frac{3 \times 400 \times 10^{-9} \times 300 \times 10^3}{1.33}) \\

θ^*_1 = sin^{-1} (0.2706) = 15.699 \;° \\

dsin θ^*_2 = \frac{mλ_2}{n} \\

λ_2 = 600 \times 10^{-9} \; m \\

θ^*_2 = sin^{-1} (\frac{3 \times 600 \times 10^{-9} \times 300 \times 10^3}{1.33}) \\

θ^*_2 = sin^{-1} (0.406) = 23.953 \;° \\

Δ θ^* = θ^*_2- θ^*_1 \\

Δ θ^* = 23.953-15.699 = 8.254 \;°"